OpenStudy (dtan5457):

Anyone wanna solve this system of equations? First one to get the right answer, gets a medal. (I just finished solving it)

OpenStudy (dtan5457):

5=y-x 4x^2=-17x+y+4

OpenStudy (nurali):

Second equation will be a quadratic if we make it all x's. y - x = 5 4x^2 + 17x - y - 4 = 0 Using the first equation, we will find y in terms of x. -y + 5 = -x y - 5 = x y = x + 5 Plug that into the second equation 4x^2 + 17x - (x + 5) - 4 =0 4x^2 + 17x - x - 9 = 0 4x^2 + 16x - 9 = 0 (2x - 1)(2x + 9) x = 0.5 or x = -4.5

Nnesha (nnesha):

5 = y-x add x both side 5 +x = y i guess this one simple instead moving y first

OpenStudy (dtan5457):

Yup, x's are correct and you'd get the y's